OpenStudy (anonymous):

∫sin(3x)cos(x)dx How would I break this up so I can evaluate?

OpenStudy (anonymous):

I got an answer last night that just confused me more.. Anyone have a clue?

OpenStudy (turingtest):

\[\cos\alpha\sin\beta=\frac12(\sin(\alpha+\beta)-\sin(\alpha-\beta))\]

OpenStudy (anonymous):

So I would end up with 1/2(sin4x)-sin(2x)??

OpenStudy (turingtest):

yes, though the 1/2 multiplies both terms

OpenStudy (anonymous):

does that make it sin2x-sinx? then I just evaluate the 2 integrals and im done?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

I can just take the 1/2 out of the integral then evaluate the 2

OpenStudy (zarkon):

yes